OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2) A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(5\quad ,&-2) \end{array} \\\quad \\ slope = m= -\cfrac{1}{5} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values}\)